Circular No. 5859

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993W IN ANONYMOUS GALAXY E. Cappellaro, European Southern Observatory, communicates: “On

Aug. 20.4 UT, I obtained a low-resolution spectrogram (range 360-700 nm, resolution 2 nm) of SN 1993W (cf. IAUC 5848) using the 3.6-m ESO telescope (+ EFOSC). The spectrum shows broad Balmer lines with a P-Cyg profile overimposed on a blue continuum. Absorption minima are measured at 429.7, 479.7, and 647.4 nm. The supernova is therefore of type II near maximum light. Provisional photometry gives B = 18.3.”

A. V. Filippenko and T. Matheson, University of California at Berkeley, report: “CCD spectrograms (range 310-1000 nm) obtained on Sept. 10 UT with the Shane 3-m reflector at Lick Observatory reveal that SN 1993W is of type II, within a few weeks of maximum brightness. The hydrogen Balmer lines, especially Hα, have prominent P-Cyg profiles. Narrow Hα emission from a faint underlying H II region and from the galaxy nucleus appears at 668 nm (z = 0.018). It should be possible to determine an accurate distance using the Expanding Photosphere Method, given enough observations.”

GX 1+4 M. H. Finger, Computer Sciences Corporation; R. B. Wilson and G.

J. Fishman, Marshall Space Flight Center, NASA; and L. Bildsten, D. Chakrabarty, and T. A. Prince, California Institute of Technology, report for the Compton Gamma Ray Observatory BATSE team: “Beginning Aug. 27, the x-ray binary pulsar GX 1+4 has dramatically brightened in the hard (> 20 keV) x-ray band. The pulsed flux is now the highest seen by BATSE in more then two years of monitoring. On Sept. 5, the barycentric pulse period was 120.567 ± 0.005 s, and the phase averaged pulsed flux (20-120 keV) is well fit with F (E) = (A/E)exp(−E/kT), where kT = 27 ± 3 and F (50 keV) = (2.1 ± 0.1) × 10^{−4} photon cm^{−2} s^{−1} keV^{−1}. The period derivative has increased steadily from 1.2 ± 0.4 s/yr as measured in 1993 March to the current rate of 3.7 ± 0.3 s/yr. The cause of the GX 1+4 outbursts is not known, but may be due to variations in the M6 III red giant companion V2116 Oph (V = 19, R = 15, k = 8; finder chart in Doxsey et al. 1977, Nature 270, 586).”

1993 September 10 Daniel W. E. Green